OpenStudy (anonymous):

solve the following equation for P P + n/4= 3n + 3

OpenStudy (gw2011):

p+n/4=3n+3 p=3n+3-n/4 p=3(n+1)-n/4 p=[12(n+1)-n]/4 p=(12n+12-n)/4 p=(11n+12)/4 p=(11/4)n+3